OpenStudy (anonymous):

I know if you have an integer exponent you flip the fraction ex)x^-2 = 1/x^2 But what if you have it in a binomial? (x+y^-1)^-1

OpenStudy (amistre64):

the rules dont change

OpenStudy (amistre64):

y^-1 = 1/y

OpenStudy (amistre64):

(yada)^-1 = 1/yada

OpenStudy (anonymous):

But you can't break up a binomial. I think you have to take the y out of the problem before you can do that

OpenStudy (amistre64):

\[(a+b)^{-5}=\frac{1}{\sum_{n=0}^{5}\ C(5,n)\ a^{5-n}\ b^{n}}\]

OpenStudy (amistre64):

what is there to break up? y is just a "number" so i dont think i understand your .... well, point

OpenStudy (anonymous):

(x+y^-1)^-1 = 1/(x+y^-1) but I don't think you can flip that y from the denominator because its part of a binomial.

OpenStudy (anonymous):

I have the answer to the problem but I don't know how you get to it. y/(xy+1)

OpenStudy (amistre64):

you dont flip the y up; "flipping" isnt the rule; it doesnt go up into the numberator; it just flipps itself over and become the reciprocal of y

OpenStudy (amistre64):

\[(x+y^{-1})^{-1}=\frac{1}{x+\frac{1}{y}}=\frac{1(y)}{(x+\frac{1}{y})(y)}=\frac{y}{xy+1}\]

OpenStudy (amistre64):

that is what going on behind the scene

OpenStudy (anonymous):

Thank you :) I see what I was doing wrong. By the way, how do you type a fraction like that?

OpenStudy (amistre64):

magic powers lol .... *[ \frac{top}{bottom} *] replace each " * " with a " \ " in order for the parser to recognize it as math code

OpenStudy (anonymous):

Ok thanks